3. If A = i 0
0 i

where i =√−1, find A^2
and A^4.
1
Expert's answer
2020-05-28T15:07:58-0400

We know that A is

"A= \begin{pmatrix}
 i & 0 \\
 0 & i
\end{pmatrix}" . Therefore, "A^2 = \begin{pmatrix}
 i & 0 \\
 0 & i
\end{pmatrix} \cdot \begin{pmatrix}
 i & 0 \\
 0 & i
\end{pmatrix} = \begin{pmatrix}
 i\cdot i+0\cdot0 & i\cdot0 +0\cdot i\\
 0\cdot i + i\cdot0 & 0\cdot0+i\cdot i
\end{pmatrix} = \begin{pmatrix}
 -1 & 0 \\
 0 & -1
\end{pmatrix}."

Next, "A^4 = A^2\cdot A^2 = \begin{pmatrix}
 -1 & 0 \\
 0 & -1
\end{pmatrix} \cdot \begin{pmatrix}
 -1 & 0 \\
 0 & -1
\end{pmatrix} = \begin{pmatrix}
 -1\cdot-1 + 0\cdot0 & -1\cdot0+0\cdot-1 \\
 0\cdot-1+-1\cdot0 & 0\cdot0+-1\cdot-1
\end{pmatrix} = \begin{pmatrix}
 1 & 0 \\
 0 &1
\end{pmatrix}"

We may also note that "A = i\cdot \begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix}," therefore "A^2 = (-1)\cdot\begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix}," "A^4 = 1\cdot \begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix} = \begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix}."